# PFML: Self-Supervised Learning of Time-Series Data Without Representation Collapse

### v2

@@ Abstract
1 Introduction
Self-supervised learning (SSL) can be described as a data-driven learning paradigm where the training process is guided by the inherent structure of the data itself. Unlike supervised learning that relies on externally provided labels, SSL exploits the intrinsic properties of the data to generate its own supervisory signal (Balestriero et al., 2023). SSL enables the model to learn rich feature representations from large amounts of unlabeled data that can be used as a starting point for downstream tasks, either as such or by fine-tuning the feature extractor to be better suited for solving some specific task (Erhan et al., 2010). Since typically there is an abundance of unlabeled data but a scarcity of labeled data, the use of SSL has been shown to reduce the need for large, manually annotated datasets (van den Oord et al., 2018; Baevski et al., 2020; Chen et al., 2020). In addition to SSL algorithms that have been developed for a single data modality, SSL algorithms that can be applied to multiple different data modalities have gained popularity in recent years (van den Oord et al., 2018; Akbari et al., 2021; Baevski et al., 2022; Wang et al., 2023). These methods and their extensions have shown great success in e.g. audio, image, and text data (van den Oord et al., 2018; Hénaff et al., 2020; Akbari et al., 2021; Baevski et al., 2022; Wang et al., 2023; Baevski et al., 2023; Yoon et al., 2023; Zhu et al., 2023; Lian et al., 2023).
However, many SSL algorithms suffer from two issues: First, SSL algorithms are usually complex, with a plethora of hyperparameters that need careful tuning for the algorithm to work properly. This hinders the ability of SSL algorithms to be applied to new data domains, where the selection of these hyperparameters is not self-evident. For example, in contrastive learning-based SSL, the selection of positive and negative samples during training is essential for the algorithm to work properly. However, deciding which samples should be assigned to positive and negative categories is not always apparent (Kalantidis et al., 2020; Robinson et al., 2021; Balestriero et al., 2023). As another example, determining the number of clusters for clustering-based SSL algorithms (such as Caron et al. (2020) and Hsu et al. (2021)) in a new data domain or task can be difficult. Examples of such domains could include, for instance, different types of medical time-series data (e.g. EEG, ECG, or EMG recordings) that come in various dataset sizes and from various recording configurations. Second, a common failure mode during SSL pre-training is representation collapse, where the model ends up outputting a constant, time-invariant feature representation. Representation collapse is very common in SSL pre-training (Hua et al., 2021; Jing et al., 2022; Balestriero et al., 2023; Garrido et al., 2023), and many SSL methods apply different countermeasures to tackle the problem (see Section 3.1).
@@ 2 Related Work
Most of the advances in SSL have focused on developing new, better-performing algorithms with some specific data modality in mind. For speech data, Baevski et al. (2020) presented an SSL algorithm where the basic idea is to mask speech embeddings and then solve a contrastive task that is defined over a quantization of the embeddings which are simultaneously learned during the pre-training task. Hsu et al. (2021) proposed that instead of solving a contrastive task, they predict cluster targets of masked embeddings. Furthermore, the SSL method by Chen et al. (2022) also uses masking of embeddings, but the authors simulate noisy speech inputs and predict pseudo-labels of the original speech from the masked embeddings.
@@
Similar to the advances in SSL for speech data, there have been significant developments in SSL for image data as well (Lee et al., 2017; Gidaris et al., 2018; Caron et al., 2018; Grill et al., 2020; Chen et al., 2020; Radford et al., 2021; He et al., 2022; Bao et al., 2022; Oquab et al., 2024). Grill et al. (2020) presented an SSL method that uses two neural networks that learn from each other’s representations of differently augmented views of the same image. He et al. (2022) proposed masked autoencoders (MAE) that try to reconstruct masked patches of input images using an asymmetric encoder-decoder architecture. The SSL algorithm by Bao et al. (2022) tokenizes images into visual tokens, followed by masking some image patches and then trying to recover the original tokens from the masked patches.
@@
SSL has also excelled in natural language processing (Devlin et al., 2019; Brown et al., 2020; Tay et al., 2023; OpenAI, 2023). Devlin et al. (2019) introduced an SSL method which obtains bidirectional feature representations from unlabeled text by conditioning on both the left and right textual context. The method by Brown et al. (2020) uses an autoregressive model which alternates dense and locally banded sparse attention patterns in their Transformer model. OpenAI (2023) proposed an expanded version of the method by Brown et al. (2020) by making the model not only larger, but also capable of handling image inputs in addition to text inputs.
@@
More recently, SSL literature has seen a growing number of work towards SSL algorithms capable of running the pre-training task on multiple different data modalities. The authors of van den Oord et al. (2018) developed an SSL approach that predicts future embeddings based on previous context using contrastive learning. They showed that their method was able to learn useful feature representations for audio, image, text, and reinforcement learning in 3D environments. The SSL method by Akbari et al. (2021) also uses contrastive learning, but their method simultaneously takes audio, video, and text data as input and creates multimodal feature representations. These features were shown to work well with multiple different downstream tasks, i.e. video action recognition, audio event classification, image classification, and text-to-video retrieval. Baevski et al. (2022) proposed data2vec, an SSL method for audio, image, and text data. In their approach, the model tries to predict masked latent features of an older version of itself that are both normalized and averaged over multiple Transformer layers. Their results in downstream tasks demonstrate the effectiveness of the method in all three data modalities. Yue et al. (2022) presented a universal framework, TS2Vec, for learning time-series representations at different semantic levels using hierarchical contrastive learning over augmented context views. Their approach enables robust contextual representations for each timestamp and allows for simple aggregation to obtain representations of arbitrary sub-sequences. They demonstrated that their method obtained state-of-the-art performance in time-series classification, forecasting, and anomaly detection tasks on multiple datasets, most of which were small in terms of the number of training samples. Notably, TS2Vec achieved these results by adding very simple classifiers after the pre-trained model, which has fixed constraints on the model architecture to model the hierarchical representations. Wang et al. (2023) proposed an SSL method that performs prediction of masked tokens in a unified manner on images, texts, and image-text pairs. Their experiments showed that their method achieves state-of-the-art performance on various vision and vision-language tasks.
@@
For modality-agnostic SSL algorithms, objective functions play a crucial role in guiding the learning process. These functions can be broadly categorized into three types: instance discrimination, clustering, and masked prediction. Instance discrimination aims to distinguish between different instances of data, thereby encouraging the model to learn unique features for each instance and enhancing the discriminative power of the learned representations. Contrastive learning methods, such as van den Oord et al. (2018); Baevski et al. (2020); He et al. (2020); Akbari et al. (2021); Pizzi et al. (2022); Yue et al. (2022), are an example of instance-discrimination-based SSL methods. Clustering, on the other hand, groups similar instances together in the feature space, fostering the model to learn common features among instances belonging to the same group. Methods like Caron et al. (2020); Asano et al. (2020); Hsu et al. (2021) are examples of clustering-based SSL methods. Lastly, masked prediction involves the task of predicting masked parts of the input data based on the unmasked parts, thereby encouraging the model to learn contextual relationships within the data. Examples of such SSL methods include Devlin et al. (2019); Wang et al. (2020); Baevski et al. (2022); He et al. (2022); Xie et al. (2022).
3 Method
3.1 Motivation
@@
One key issue with many SSL methods is the problem of representation collapse, where the model outputs a constant, input-invariant feature representation, leading to a trivial solution of the pre-training task (Jing et al., 2022; Balestriero et al., 2023). This considerably slows down the development process for novel data domains and/or tasks due to the necessity of operating in uncertainty, when it is not clear whether the representation collapse is caused by an ill-posed task or by the SSL algorithm. To avoid this, SSL methods have taken several different countermeasures: Baevski et al. (2020) use the same target representations in their contrastive learning task in a dual manner, i.e. both as a positive and a negative example. Grill et al. (2020) both add an additional predictor to their training regime and use a moving average of their so-called online neural network to avoid representation collapse. Bardes et al. (2022) add a regularization term to their loss function that both maintains variance of the embeddings and decorrelates each pair of variables. In data2vec (Baevski et al., 2022), the authors tackle representation collapse by carefully selecting their model hyperparameters and promoting target representation variance through feature normalization. Also, in the code implementation of data2vec [^1] https://github.com/facebookresearch/fairseq/tree/main/examples/data2vec, pre-training is stopped if the variance of either model predictions or training targets falls below a predefined threshold.
@@
Intuitively, given a trivial task, the model does not learn useful feature representations during pre-training. In contrast, if the learning objective is too complicated, the model fails to converge to a useful solution. For time-series data, i.e. a waveform (e.g. audio) or a set of waveforms (e.g. multi-channel EEG), trying to reconstruct masked parts of the input signal given the unmasked parts of the signal (as in e.g. MAE (He et al., 2022)) is a very complex task. This is due to the fact that a time-series signal can have large temporal variation even between short periods of time. While joint learning of a priori unspecified latent representations and their prediction allows discarding of this irrelevant variation (as in, e.g., van den Oord et al. (2018) or Baevski et al. (2020)), the problem requires learning algorithms that become susceptible to representation collapse and/or may require careful tuning of the training process. We hypothesize that for SSL pre-training with time-series data, a model would learn more useful features for downstream tasks if the complex setting of MAE would be alleviated slightly. Hence, we propose Prediction of Functionals from Masked Latents (PFML), a novel SSL algorithm for time-series data. Our method builds on the concept of MAE and reduces the complexity of the pre-training task of MAE in two ways:
@@
    Instead of masking the input signal directly, PFML borrows the idea of e.g. wav2vec 2.0 (Baevski et al., 2020) and data2vec (Baevski et al., 2022) and masks the embeddings created by the encoder model.
Regarding point (1), by making the model predict statistical functionals of masked latent features instead of predicting the input signal $\mathbf{x}$ itself, we relieve the model from the complex task of modelling the high-dimensional distribution of $\mathbf{x}$ in detail. We validate this argument of generating better features for downstream tasks by reducing the computational complexity of the pre-training task in Section 4, where we compare our proposed method against MAE. In theory, the set of statistical functionals can be chosen so that the desired and deterministically calculated statistical properties of the data, and thereby their variance, are preserved in the target features. Furthermore, regarding point (2), we show in our experiments in Section 4 that it is more beneficial during pre-training to mask the latent features instead of masking the input directly. This further alleviates the complexity of the learning task in particular for the encoder module.
3.2 Prediction of Functionals from Masked Latents
Figure 1: An overview of the PFML pre-training pipeline.
@@ 3.2 Prediction of Functionals from Masked Latents
In PFML, predicting only one or a few functionals of a framed signal can be a trivial task, and will most probably lead to learning feature representations that are not very useful for downstream tasks. However, as the number of functionals that each describe some property of the framed signal grows, a more accurate description of the signal can be obtained (see e.g. McDermott & Simoncelli (2011) for representing perceptual properties of sound textures with functionals). Therefore, as the number of different functionals grows, the PFML algorithm is getting closer to predicting all of the nuances of the input signal.
@@
Under these assumptions, as the model is trying to predict the computed functionals $\mathbf{f}_{n}$ given the embeddings $\mathbf{z}_{n}$, good model predictions $\mathbf{y}_{n}$ that lead to low prediction loss values also inherently contain variance. On the contrary, if $\mathbf{y}_{n}$ were to contain zero variance across the frames while $\mathbf{f}_{n}$ contains variance, the prediction loss would be high. Consequently, PFML pre-training does not converge to collapsed feature representations, as long as Assumptions 1 and 2 hold true. For a more detailed formulation, see Appendix A. Empirical results (see Section 5) support this theoretical claim, showing that PFML maintains variance in predictions across various datasets.
@@
where sgn denotes the sign function (Rabiner & Schafer, 2007). The ACF for a signal $\mathbf{x}$ at lag $\tau$ is defined as
@@
where $\tau<N$, $\mu$ is the mean of $\mathbf{x}$, and $\sigma^{2}$ is the variance of $\mathbf{x}$ (Rabiner & Schafer, 2007). Note that Equation 2 returns a vector of measurements when applied to all lags $\tau<N$.
@@
Note that the PFML pre-training process is not restricted to any specific type of neural networks. In the present study, we used convolutional neural networks (CNNs) as our encoder model, and $T$ Transformer encoder blocks as the temporal model. However, any type of encoder could be used for PFML, as long as the encoder can convert time-series data into a sequence of embeddings. Furthermore, other temporal models, such as conformer-based models (Gulati et al., 2020) or bidirectional recurrent neural networks (Hochreiter & Schmidhuber, 1997; Cho et al., 2014), could also be used for PFML, as long as the model is able to take contextual information into account.
4 Experiments
@@
We evaluate our PFML method using three different datasets of time-series data with complex classification tasks: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. For each dataset, we first run SSL pre-training with unlabeled data using PFML, after which we fine-tune our models for downstream classification tasks using labeled data. We compare PFML against four different baselines: MAE (He et al., 2022), data2vec (Baevski et al., 2022), TS2Vec (Yue et al., 2022), and not using pre-training at all. We selected MAE for our experiments since it is conceptually very similar to PFML, and we chose data2vec since it is the current state-of-the-art data-modality-agnostic SSL method. We also included the conceptually different TS2Vec in the present experiments due to its reported state-of-the-art performance on multiple different datasets of single- and multi-channel time-series data. In order to make the prediction of functionals directly comparable with predicting the input signal, we use a slightly modified version of MAE where we mask embeddings instead of masking inputs.
@@
This section is organized as follows: First, Section 4.1 gives general-level information on the pre-training and fine-tuning experiments regarding PFML, MAE, and data2vec methods that utilize identical neural network architecture for the end-use system. This is followed by Sections 4.2, 4.3, and 4.4, which give modality-specific experimental details for IMU, speech, and EEG data, respectively. Finally, Section 4.5 explains the key differences between TS2Vec-related experiments and the experiments described in Section 4.1.
4.1 PFML, MAE, and data2vec Experiments

### v3

@@ Abstract
Index Terms:
EEG data, embedding masking, human activity recognition, multi-sensor inertial measurement unit data, representation collapse, self-supervised learning, sleep stage classification, speech emotion recognition, statistical functionals, time-series data.
[…]
=-21pt
I Introduction
Self-supervised learning (SSL) can be described as a data-driven learning paradigm where the training process is guided by the inherent structure of the data itself. Unlike supervised learning that relies on externally provided labels, SSL exploits the intrinsic properties of the data to generate its own supervisory signal [1, 2]. SSL enables the model to learn rich feature representations from large amounts of unlabeled data that can be used as a starting point for downstream tasks, either as such or by fine-tuning the feature extractor to be better suited for solving some specific task [2, 3]. Since typically there is an abundance of unlabeled data but a scarcity of labeled data, the use of SSL has been shown to reduce the need for large, manually annotated datasets [4, 5, 6]. In addition to SSL algorithms that have been developed for a single data modality, SSL algorithms that can be applied to multiple different data modalities have gained popularity in recent years [2, 4, 7, 8, 9]. These methods and their extensions have shown great success in e.g. audio, image, and text data [4, 10, 7, 8, 9, 11, 12, 13, 14].
However, many SSL algorithms suffer from two issues: First, SSL algorithms are usually complex, with a plethora of hyperparameters that need careful tuning for the algorithm to work properly. This hinders the ability of SSL algorithms to be applied to new data domains, where the selection of these hyperparameters is not self-evident. For example, in contrastive learning-based SSL, the selection of positive and negative samples during training is essential for the algorithm to work properly. However, deciding which samples should be assigned to positive and negative categories is not always apparent [15, 16, 1]. As another example, determining the number of clusters for clustering-based SSL algorithms (such as Caron et al. [17] and Hsu et al. [18]) in a new data domain or task can be difficult. Examples of such domains could include, for instance, different types of medical time-series data (e.g. electroencephalography (EEG), electrocardiography (ECG), or electromyography (EMG) recordings) that come in various dataset sizes and from various recording configurations. Second, a common failure mode during SSL pre-training is representation collapse, where the model ends up outputting a constant, time-invariant feature representation. Representation collapse is very common in SSL pre-training [19, 20, 1, 21], and many SSL methods apply different countermeasures to tackle the problem (see Section III-A).
@@ II Related Work
Most of the advances in SSL have focused on developing new, better-performing algorithms with some specific data modality in mind. For speech data, Baevski et al. [5] presented an SSL algorithm where the basic idea is to mask speech embeddings and then solve a contrastive task that is defined over a quantization of the embeddings which are simultaneously learned during the pre-training task. Hsu et al. [18] proposed that instead of solving a contrastive task, they predict cluster targets of masked embeddings. Furthermore, the SSL method by Chen et al. [22] also uses masking of embeddings, but the authors simulate noisy speech inputs and predict pseudo-labels of the original speech from the masked embeddings.
@@
Similar to the advances in SSL for speech data, there have been significant developments in SSL for image data as well [23, 24, 25, 26, 6, 27, 28, 29, 30]. Grill et al. [26] presented an SSL method that uses two neural networks that learn from each other’s representations of differently augmented views of the same image. He et al. [28] proposed masked autoencoders (MAE) that try to reconstruct masked patches of input images using an asymmetric encoder-decoder architecture. The SSL algorithm by Bao et al. [29] tokenizes images into visual tokens, followed by masking some image patches and then trying to recover the original tokens from the masked patches.
@@
SSL has also excelled in natural language processing [31, 32, 33, 34]. Devlin et al. [31] introduced an SSL method which obtains bidirectional feature representations from unlabeled text by conditioning on both the left and right textual context. The method by Brown et al. [32] uses an autoregressive model which alternates dense and locally banded sparse attention patterns in their Transformer model. OpenAI [34] proposed an expanded version of the method by Brown et al. [32] by making the model not only larger, but also capable of handling image inputs in addition to text inputs.
@@
More recently, SSL literature has seen a growing number of work towards SSL algorithms capable of running the pre-training task on multiple different data modalities. The authors of van den Oord et al. [4] developed an SSL approach that predicts future embeddings based on previous context using contrastive learning. They showed that their method was able to learn useful feature representations for audio, image, text, and reinforcement learning in 3D environments. The SSL method by Akbari et al. [7] also uses contrastive learning, but their method simultaneously takes audio, video, and text data as input and creates multimodal feature representations. These features were shown to work well with multiple different downstream tasks, i.e. video action recognition, audio event classification, image classification, and text-to-video retrieval. Baevski et al. [8] proposed data2vec, an SSL method for audio, image, and text data. In their approach, the model tries to predict masked latent features of an older version of itself that are both normalized and averaged over multiple Transformer layers. Their results in downstream tasks demonstrate the effectiveness of the method in all three data modalities. Yue et al. [35] presented a universal framework, TS2Vec, for learning time-series representations at different semantic levels using hierarchical contrastive learning over augmented context views. Their approach enables robust contextual representations for each timestamp and allows for simple aggregation to obtain representations of arbitrary sub-sequences. They demonstrated that their method obtained state-of-the-art performance in time-series classification, forecasting, and anomaly detection tasks on multiple datasets, most of which were small in terms of the number of training samples. Notably, TS2Vec achieved these results by adding very simple classifiers after the pre-trained model, which has fixed constraints on the model architecture to model the hierarchical representations. Wang et al. [9] proposed an SSL method that performs prediction of masked tokens in a unified manner on images, texts, and image-text pairs. Their experiments showed that their method achieves state-of-the-art performance on various vision and vision-language tasks.
@@
For modality agnostic SSL algorithms, objective functions play a crucial role in guiding the learning process. These functions can be broadly categorized into three types: instance discrimination, clustering, and masked prediction. Instance discrimination aims to distinguish between different instances of data, thereby encouraging the model to learn unique features for each instance and enhancing the discriminative power of the learned representations. Contrastive learning methods, such as [4, 5, 36, 7, 37, 35], are an example of instance discrimination-based SSL methods. Clustering, on the other hand, groups similar instances together in the feature space, fostering the model to learn common features among instances belonging to the same group. Methods like [17, 38, 18] are examples of clustering-based SSL methods. Lastly, masked prediction involves the task of predicting masked parts of the input data based on the unmasked parts, thereby encouraging the model to learn contextual relationships within the data. Examples of such SSL methods include [31, 39, 8, 28, 40].
Figure 1: An overview of the PFML pre-training pipeline. Note that in the figure, the input signal has only a single channel, whereas PFML can also be applied to multi-channel time-series data.
III Method
III-A Motivation
@@
One key issue with many SSL methods is the problem of representation collapse, where the model outputs a constant, input-invariant feature representation, leading to a trivial solution of the pre-training task [20, 1]. This considerably slows down the development process for novel data domains and/or tasks due to the necessity of operating in uncertainty, when it is not clear whether the representation collapse is caused by an ill-posed task or by the SSL algorithm. To avoid this, SSL methods have taken several different countermeasures: Baevski et al. [5] use the same target representations in their contrastive learning task in a dual manner, i.e. both as a positive and a negative example. Grill et al. [26] both add an additional predictor to their training regime and use a moving average of their so-called online neural network to avoid representation collapse. Bardes et al. [41] add a regularization term to their loss function that both maintains variance of the embeddings and decorrelates each pair of variables. In data2vec [8], the authors tackle representation collapse by carefully selecting their model hyperparameters and promoting target representation variance through feature normalization. Also, in the code implementation of data2vec [^1] https://github.com/facebookresearch/fairseq/tree/main/examples/data2vec, pre-training is stopped if the variance of either model predictions or training targets falls below a predefined threshold.
@@
Intuitively, given a trivial task, the model does not learn useful feature representations during pre-training. In contrast, if the learning objective is too complicated, the model fails to converge to a useful solution. For time-series data, i.e. a waveform (e.g. audio) or a set of waveforms (e.g. multi-channel EEG), trying to reconstruct masked parts of the input signal given the unmasked parts of the signal (as in e.g. MAE [28]) is a very complex task. This is due to the fact that a time-series signal can have large temporal variation even between short periods of time. While joint learning of a priori unspecified latent representations and their prediction allows discarding of this irrelevant variation (as in, e.g., [4] or [5]), the problem requires learning algorithms that become susceptible to representation collapse and/or may require careful tuning of the training process. We hypothesize that for SSL pre-training with time-series data, a model would learn more useful features for downstream tasks if the complex setting of MAE would be alleviated slightly. Hence, we propose PFML, a novel SSL algorithm for time-series data. Our method builds on the concept of MAE and reduces the complexity of the pre-training task of MAE in two ways:
@@
    Instead of masking the input signal directly, PFML borrows the idea of e.g. wav2vec 2.0 [5] and data2vec [8] and masks the embeddings created by the encoder model.
Regarding point (1), by making the model predict statistical functionals of masked latent features instead of predicting the input signal $\mathbf{x}$ itself, we relieve the model from the complex task of modelling the high-dimensional distribution of $\mathbf{x}$ in detail. We validate this argument of generating better features for downstream tasks by reducing the computational complexity of the pre-training task in Section IV, where we compare our proposed method against MAE. In theory, the set of statistical functionals can be chosen so that the desired and deterministically calculated statistical properties of the data, and thereby their variance, are preserved in the target features. Furthermore, regarding point (2), we show in our experiments in Section IV that it is more beneficial during pre-training to mask the latent features instead of masking the input directly. This further alleviates the complexity of the learning task in particular for the encoder module.
III-B Prediction of Functionals from Masked Latents
@@
In PFML, predicting only one or a few functionals of a framed signal can be a trivial task, and will most probably lead to learning feature representations that are not very useful for downstream tasks. However, as the number of functionals that each describe some property of the framed signal grows, a more accurate description of the signal can be obtained (see e.g. McDermott & Simoncelli [42] for representing perceptual properties of sound textures with functionals). Therefore, as the number of different functionals grows, the PFML algorithm is getting closer to predicting all of the nuances of the input signal.
@@
Under these assumptions, as the model is trying to predict the computed functionals $\mathbf{f}_{n}$ given the embeddings $\mathbf{z}_{n}$, good model predictions $\mathbf{y}_{n}$ that lead to low prediction loss values also inherently contain variance. On the contrary, if $\mathbf{y}_{n}$ were to contain zero variance across the frames while $\mathbf{f}_{n}$ contains variance, the prediction loss would be high. Consequently, PFML pre-training does not converge to collapsed feature representations, as long as Assumptions 1 and 2 hold true. For a more detailed formulation, see Appendix A in the supplementary material. Empirical results (see Section V) support this theoretical claim, showing that PFML maintains variance in predictions across various datasets.
@@
where sgn denotes the sign function [43]. The ACF for a signal $\mathbf{x}$ at lag $\tau$ is defined as
@@
where $\tau<N$, $\mu$ is the mean of $\mathbf{x}$, and $\sigma^{2}$ is the variance of $\mathbf{x}$ [43]. Note that Equation 2 returns a vector of measurements when applied to all lags $\tau<N$.
@@
Note that the PFML pre-training process is not restricted to any specific type of neural networks. In the present study, we used convolutional neural networks (CNNs) as our encoder model, and $T$ Transformer encoder blocks as the temporal model. However, any type of encoder could be used for PFML, as long as the encoder can convert time-series data into a sequence of embeddings. Furthermore, other temporal models, such as conformer-based models [44] or bidirectional recurrent neural networks [45, 46], could also be used for PFML, as long as the model is able to take contextual information into account.
IV Experiments
@@
We evaluate our PFML method using three different datasets of time-series data with complex classification tasks: infant posture and movement classification from multi-sensor IMU data, emotion recognition from speech data, and sleep stage classification from EEG data. For each dataset, we first run SSL pre-training with unlabeled data using PFML, after which we fine-tune our models for downstream classification tasks using labeled data. We compare PFML against four different baselines: MAE [28], data2vec [8], TS2Vec [35], and not using pre-training at all. We selected MAE for our experiments since it is conceptually very similar to PFML, and we chose data2vec since it is the current state-of-the-art data-modality-agnostic SSL method. We also included the conceptually different TS2Vec in the present experiments due to its reported state-of-the-art performance on multiple different datasets of single- and multi-channel time-series data. In order to make the prediction of functionals directly comparable with predicting the input signal, we use a slightly modified version of MAE where we mask embeddings instead of masking inputs.
@@
This section is organized as follows: First, Section IV-A gives general-level information on the pre-training and fine-tuning experiments regarding PFML, MAE, and data2vec methods that utilize identical neural network architecture for the end-use system. This is followed by Sections IV-B, IV-C, and IV-D, which give modality-specific experimental details for IMU, speech, and EEG data, respectively. Finally, Section IV-E explains the key differences between TS2Vec-related experiments and the experiments described in Section IV-A.
IV-A PFML, MAE, and data2vec Experiments